OpenStudy (anonymous):

Find the center of the ellipse (x-4)^(2)+(y-1)^(2)/(16)=1

OpenStudy (watchmath):

The center simply the number inside your parenthesis, i.e. (4,1).

OpenStudy (anonymous):

There are many numbers inside the parantheses

OpenStudy (anonymous):

watchmath, did you do the ring problem, or were you just being challenging?

OpenStudy (watchmath):

I did it once, but I am not sure if I can remember the solution :).

OpenStudy (anonymous):

Were you saying that (4,1) is the answer?

OpenStudy (watchmath):

yes :)

OpenStudy (anonymous):

thank you

OpenStudy (watchmath):

Whenever you have \[\frac{(x-h)^2}{a^2}+\frac{(x-k)^2}{b^2}=1\] then the center is \((h,k)\).

OpenStudy (anonymous):

then wouldn't it be (-4,-1)? or no, because of the squares?

OpenStudy (watchmath):

Compare to \[\frac{(x-4)^2}{1^2}+\frac{(y-1)^2}{16}=1\] What is the \((h,k)\) here?

OpenStudy (anonymous):

But it wouldn't be negative coordinates, it would be positive?

OpenStudy (watchmath):

well if you compare (x-h) to (x-4) then the h is 4 in this case. Similarly if you compare (y-k) and (y-1) then k=1 here. So (4,1) is the center.

OpenStudy (anonymous):

okay thanks